OpenStudy (kmeezy):

Can someone please help me find the cubic equation of a graph with the coordinates: (0,1) (1,2.5) (3,1) (4.5,4)

Directrix (directrix):

If you have four distinct points in the xy-plane, and no two x-coordinates are equal, then there is a unique cubic equation of the form y = ax³ + bx² + cx + d that passes through the four points.

Directrix (directrix):

For (0,1), x = 0 and y = 1. y = ax³ + bx² + cx + d becomes for that point: 1 = a*0³ + b*0² + c*0 + d 1 = 0 + 0 + 0 + d d = 1 That leaves a, b, and c to find.

Directrix (directrix):

For (1, 2.5), x = 1 and y = 2.5 For that point, y = ax³ + bx² + cx + d becomes: 2.5 = a*1³ + b*1² + c*1 + d d is known to be 1 2.5 = a + b + c + 1 1.5 = a + b + c

Directrix (directrix):

Go through this procedure twice more, once for (3,1) and once for (4.5,4). @kmeezy

OpenStudy (kmeezy):

I did and I got 0= 9a+3b+c and 3= 91.125a+20.25b+4.5c but i do not know how to proceed after this Also I have attached the graph over here

OpenStudy (kmeezy):

@Directrix

Directrix (directrix):

For the third equation, I got this; 0= 27*a + 9*b + 3*c

OpenStudy (kmeezy):

Yeah i simplified it, if that is right?

Directrix (directrix):

We have 3 equations in 3 unknowns so we have to solve that system for a, b, and c. 1.5 = a + b + c 0 = 9*a+3*b + c 3= 91.125*a + 20.25*b + 4.5*c

Directrix (directrix):

I think c is the easiest variable to eliminate. That would suppress the system to a 2 by 2 system.

Directrix (directrix):

Are you supposed to show all the work on this?

OpenStudy (kmeezy):

Maybe this is very complicated because I do not know the accurate coordinates, did you see the graph? No actually i have to plot the second derivative of this graph.

Directrix (directrix):

>>Can someone please help me find the cubic equation of a graph with the coordinates: (0,1) (1,2.5) (3,1) (4.5,4) I worked on the posted question. If you need to plot the curve of the second derivative, then I think you should crank out the values of a, b, and c in the 3 by 3 system. Then, you can write the exact equation of the function. After that, take its derivate and draw the graph of that.

OpenStudy (kmeezy):

I have figured out the equation it is : 1/3x^3-2x^2+3x+1 Now the derivative part is easy Thank you so much for helping!

Directrix (directrix):

You are welcome.